OpenStudy (anonymous):

I really need some help, please! Write the equation of a circle with center (3,-9) and radius r=4

OpenStudy (megannicole51):

equation of a circle formula is (x-h)^2+(y-k)^2=r^2 so just plug in ur x and y values, then since u have your radius...just square it:)

terenzreignz (terenzreignz):

We have a pattern, or formula, if you will, for this type of problem. If you have a radius r and a center (h,k) then the equation of the circle looks like this \[\Large (x-h)^2 + (y-k)^2 = r^2\] In your case, h = 3 k = -9 and of course, r = 4 Just plug in and be done :)